# Stability analysis of very high order minimization-based and Taylor-based embedded boundary treatments of discontinuous Galerkin for hyperbolic equations

### v1

@@ 5.3 ROD treatment and explicit time integration
Following the same configuration considered above for the SB treatment. The analysis for the ROD boundary treatment can be performed thanks to the reformulations carried out in section 4, which allows one to write easily the ROD correction in matrix form.
@@
and the stability regions are given in figure 5. Contrary to the SB correction, the ROD-E method with polynomial degree $p=1,2,3$ shows stable areas for all values of $d\in[-1,0]$, with no constraint on the CFL. Instead, for $d\in[0,1]$, the method becomes unstable after a certain threshold $d>d_{max}$, which depends on the polynomial degree used to build the scheme, with the same $d_{max}$ observed for the SB method. Although the ROD-E method shows better stability properties for $p=1,2,3$ and distance $d\in[-1,0]$ than the SB method, for higher order polynomials $p=4,5,6$, also the ROD-E method becomes unconditionally unstable after a certain threshold. Contrary to the SB method, the ROD-E method never shows a constraint on the CFL number, it is either stable under a classical CFL condition or it becomes unconditionally unstable after a certain distance threshold.
As done in section 3, we can analytically compute the eigenvalues of the ROD-E boundary discretized operator. The analytical eigenvalues of the $\mathbb{P}^{1}$ method are
@@
Also in this case, the scheme is stable in the semi-discrete sense when $d<2/3\approx 0.66$, which is the distance threshold after which the scheme becomes unconditionally unstable.
@@
5.4 Shifted boundary treatment and implicit time integration
As we did in our previous work on the subject [ciallella2025stability], the possibility of stabilizing the method through implicit time integration is investigated for very high order methods. For this reason we consider the best time integration method in terms of stability, which is implicit Euler. Contrary to [ciallella2025stability], where the analysis was done only for $p=1,2,3$, here it is investigated whether the very high order embedded treatment can also benefit from implicit time integration.
If we consider the implicit Euler time integration of the discontinuous Galerkin method with the SB polynomial correction, the global system reads
$\begin{split}M_{1}\frac{\mathbf{u}^{n+1}_{1}-\mathbf{u}^{n}_{1}}{\Delta t}&=K^{s}_{1}\mathbf{u}^{n+1}_{1}-(K^{R}_{1}\mathbf{u}^{n+1}_{1}-K^{SB}_{1}\mathbf{u}^{n+1}_{1}),\\ M_{2}\frac{\mathbf{u}^{n+1}_{2}-\mathbf{u}^{n}_{2}}{\Delta t}&=K^{s}_{2}\mathbf{u}^{n+1}_{2}-(K^{R}_{2}\mathbf{u}^{n+1}_{2}-K^{L}_{2}\mathbf{u}^{n+1}_{1}),\end{split}$ (57)
In this case, the system reduces to finding the eigenvalues of the following amplification matrix:
$\begin{bmatrix}\mathbf{u}^{n+1}_{1}\\ \mathbf{u}^{n+1}_{2}\end{bmatrix}=\left[I-\Delta t\mathbf{M}^{-1}\mathbf{K}\right]^{-1}\begin{bmatrix}\mathbf{u}^{n}_{1}\\ \mathbf{u}^{n}_{2}\end{bmatrix}$ (58)
where the global matrices are those defined in equation (55).
The stability regions for a $\text{CFL}\in[0,1]$ are shown in figure 7. It is interesting to notice that the implicit time integration is able to improve the stability regions of the SB method, at least for values of $d\in[-1,0]$ and $p=1,2,3,4$. However, when going very high order for $p=5,6$, instabilities start to arise even for implicit time integration and a CFL greater than a certain threshold is necessary to stabilize. In particular, stability for higher order cases imposes a lower bound on the CFL condition: $\text{CFL}\geq 0.7$ for $\mathbb{P}^{5}$, and $\text{CFL}\geq 2.5$ for $\mathbb{P}^{6}$. This is shown in figure 8, where the stability regions for $\text{CFL}\in[0,10]$ are plotted.
Figure 7: Stability region: SB correction of the homogeneous Dirichlet condition for the implicit discontinuous Galerkin method and $\text{CFL}\in[0,1]$. Green areas are stable, while red areas are unstable.
Figure 8: Stability region: SB correction of the homogeneous Dirichlet condition for the implicit discontinuous Galerkin method and $\text{CFL}\in[0,10]$. Green areas are stable, while red areas are unstable.
@@ 5.5 ROD treatment and implicit time integration
Similar as above the same thing can be done to study the effect of implicit time integration when dealing with ROD treatment. If we consider the implicit Euler time integration of the discontinuous Galerkin method with the ROD (either ROD-E or ROD-$L^{2}$) polynomial correction, the global system reads
@@ 6 Numerical experiments
For the numerical experiments, the linear advection equation with source term is set up to preserve a stationary manufactured solution given by the following exact solution
$u_{ex}(x)=0.1\sin(\pi x),$
which is also taken as initial condition. The source term is computed as
$\partial_{t}u_{ex}(x)+\partial_{x}u_{ex}(x)=s(x),\quad\text{that gives}\quad s(x)=0.1\pi\cos(\pi x).$
@@
The computational domain is $\Omega=[0,2]$ with a Dirichlet boundary condition imposed at the left boundary. No conditions are imposed at the right boundary due to the pure upwind flux. A set of uniform meshes with $N_{e}=20,40,80,160$ elements is considered for $\mathbb{P}^{p}$ with $p=1,2,3$. While another set of uniform meshes with $N_{e}=5,10,20,40$ elements is considered for $\mathbb{P}^{p}$ with $p=4,5,6$.
@@
The SB and ROD (both ROD-E and ROD-$L^{2}$) polynomial corrections are imposed considering different distance values $d$ of the boundary from the left interface of the first cell. When $d=0$, the boundary coincides with the left interface of the first cell, and therefore the correction is not applied. This is the case of standard fitted boundary conditions. Since all studied methods present more instabilities when $d\in[0,\Delta x]$, the distances considered in the numerical experiments will be taken as $d\in[-\Delta x,0]$, meaning with the real boundary point collocated outside the first mesh element. For simplicity, in convergence tables, the distance is reported in non-dimensional form, i.e. $d/\Delta x$. Stability results obtained through the analysis are perfectly corroborated by the numerical results for all configurations of $d$, and for both explicit and implicit time integration.
In table 1, we present the numerical results obtained with the SB correction and explicit time integration. We consider the greater distance that makes the scheme stable under the associated maximum CFL value. As also mentioned above, for $d<0$, it is possible reach the maximum distance of d=1d=-1 for polynomials $p=1,2,3,4$ with a strict CFL condition that becomes stricter as $p$ increases. For $p=5,6$, the scheme is unconditionally unstable for d=1d=-1, and the maximum distance value for such high order polynomials is much lower: d=0.25d=-0.25 for $\mathbb{P}^{5}$, and d=0.07d=-0.07 for $\mathbb{P}^{6}$.
Table 1: Linear advection equation: $L_{2}$ errors $\|u-u_{ex}\|_{2}$ and estimated order of accuracy (EOA) with SB polynomial correction and explicit time integration for polynomials $\mathbb{P}^{p}$. CFL and $d$ are chosen to consider the worst case scenario, i.e. the furthest possible $d\in[-1,0]$ and the associated maximum CFL.
$\mathbb{P}^{1}$, CFL=6E-1, d=1.00d=-1.00 $\mathbb{P}^{2}$, CFL=2E-1, d=1.00d=-1.00 $\mathbb{P}^{3}$, CFL=5E-2, d=1.00d=-1.00
$N_{e}$, $L_{2}$ error EOA $L_{2}$ error EOA $L_{2}$ error EOA
20 5.98E-04 2.68E-03 2.59E-05
40 1.27E-04 2.23 3.37E-04 3.00 8.12E-07 4.99
80 3.02E-05 2.07 4.22E-05 3.00 2.54E-08 5.00
160 7.45E-06 2.01 5.28E-06 3.00 7.65E-10 5.04
$\mathbb{P}^{4}$, CFL=8E-3, d=1.00d=-1.00 $\mathbb{P}^{5}$, CFL=8E-2, d=0.25d=-0.25 $\mathbb{P}^{6}$, CFL=3E-1, d=0.07d=-0.07
$N_{e}$, $L_{2}$ error EOA $L_{2}$ error EOA $L_{2}$ error EOA
5 2.58E-02 4.80E-05 4.95E-07
10 8.89E-04 4.86 3.96E-07 6.92 4.52E-09 6.77
20 2.84E-05 4.96 3.13E-09 6.98 3.67E-11 6.94
40 8.94E-07 4.99 2.46E-11 6.99 2.91E-13 6.97
@@
In table 2, we present the numerical results obtained with the ROD-E correction and explicit time integration. Again, we consider the greater distance that makes the scheme stable under the associated maximum CFL value. As also mentioned above, for $d<0$, it is possible reach the maximum distance of d=1d=-1 for polynomials $p=1,2,3$ with no condition on the CFL. For $p=4,5,6$, the scheme is unconditionally unstable for d=1d=-1, and the maximum distance value for such polynomials is much lower: d=0.10d=-0.10 for $\mathbb{P}^{4}$, d=0.04d=-0.04 for $\mathbb{P}^{5}$, and d=0.015d=-0.015 for $\mathbb{P}^{6}$. Contrary to the SB method, there are no cases where it is possible to stabilize the ROD-E method by lowering the CFL. ROD-E is either stable under a classical CFL condition or unconditionally unstable, and this is what happens for higher order polynomials after a certain distance threshold.
@@
In table 3, we present the numerical results obtained with the ROD-$L^{2}$ correction and explicit time integration. Again, we consider the greater distance that makes the scheme stable under the associated maximum CFL value. As also mentioned above, for $d<0$, it is possible reach the maximum distance of d=1d=-1 for polynomials $p=1,2,3,4$ with no condition on the CFL, which means that we have the method $p=4$ that is stable contrary to ROD-E. For $p=5,6$, the scheme is unconditionally unstable for d=1d=-1, and the maximum distance value for such polynomials is much lower: d=0.25d=-0.25 for $\mathbb{P}^{5}$, and d=0.05d=-0.05 for $\mathbb{P}^{6}$. Also in this case, there are no cases where it is possible to stabilize the ROD-$L^{2}$ method by lowering the CFL. ROD-$L^{2}$ is either stable under a classical CFL condition or unconditionally unstable, and this is what happens for higher order polynomials after a certain distance threshold, similarly to what was observed for ROD-E.
@@
In table 4, we present the numerical results obtained with the SB correction and implicit time integration. For implicit time integration, it is always possible to obtain a stable scheme for d=1d=-1. However, as it was studied above, for some cases it necessary to have a CFL higher than a certain threshold. For the SB method with $p=1,2,3,4$ there are no lower bound conditions on the CFL, and it is therefore possible to use any CFL value and the scheme will stays stable. However, for $p=5$ it is necessary to have $\text{CFL}\geq 0.6$, while for $p=6$ we need a $\text{CFL}\geq 2.5$. This means that even with implicit time integration, the higher order methods still have some parasitic modes that need higher CFL values to be damped out.
Table 4: Linear advection equation: $L_{2}$ errors $\|u-u_{ex}\|_{2}$ and estimated order of accuracy (EOA) with SB polynomial correction and implicit time integration for polynomials $\mathbb{P}^{p}$. It is always considered the case where d=1.0d=-1.0 with the smallest CFL $\geq 1$ value to make the scheme stable.
$\mathbb{P}^{1}$, CFL=1.0 $\mathbb{P}^{2}$, CFL=1.0 $\mathbb{P}^{3}$, CFL=1.0
$N_{e}$, $L_{2}$ error EOA $L_{2}$ error EOA $L_{2}$ error EOA
20 5.98E-04 2.68E-03 2.59E-05
40 1.27E-04 2.23 3.37E-04 3.00 8.12E-07 4.99
80 3.02E-05 2.07 4.22E-05 3.00 2.54E-08 5.00
160 7.45E-06 2.01 5.28E-06 3.00 7.65E-10 5.04
$\mathbb{P}^{4}$, CFL=1.0 $\mathbb{P}^{5}$, CFL=1.0 $\mathbb{P}^{6}$, CFL=2.5
$N_{e}$, $L_{2}$ error EOA $L_{2}$ error EOA $L_{2}$ error EOA
5 2.58E-02 3.48E-03 2.04E-03
10 8.89E-04 4.86 2.85E-05 6.93 1.76E-05 6.85
20 2.84E-05 4.96 2.26E-07 6.98 1.41E-07 6.96
40 8.94E-07 4.99 1.77E-09 6.99 1.18E-09 6.90
In table 5, we present the numerical results obtained with the ROD-E correction and implicit time integration. As for the SB method, for implicit time integration, it is always possible to obtain a stable scheme for d=1d=-1 but for some cases the considered method has a lower bound condition on the CFL. For the ROD-E method with $p=1,2,3$ there a no conditions on the CFL, and it is therefore possible to use any CFL value and the scheme will stays stable. This was expected since these schemes were also stable under standard CFL condition with explicit time integration. In this case, it is interesting to see if coupling higher order methods with implicit time integration can result in stable schemes. However, for $p=4,5,6$ a lower bound on the CFL is necessary to ensure stability: for $p=4$ it is necessary to have $\text{CFL}\geq 3$, for $p=5$ we need $\text{CFL}\geq 6$, while for $p=6$ we have $\text{CFL}\geq 9$.
@@
In table 6, we present the numerical results obtained with the ROD-$L^{2}$ correction and implicit time integration. Also in this case, it is always possible to obtain a stable scheme for d=1d=-1 but for some cases the considered method has a lower bound condition on the CFL. For the ROD-$L^{2}$ method with $p=1,2,3,4$ there a no conditions on the CFL, and it is therefore possible to use any CFL value and the scheme will stays stable. This was expected since these schemes were also stable under standard CFL condition with explicit time integration. Once again, we study here if coupling higher order methods with implicit time integration can improve the stability for $p=5,6$. We show that for $p=5$ it is necessary to have $\text{CFL}\geq 0.7$, while for $p=6$ we need $\text{CFL}\geq 2$. Although less strong than in the ROD-E case, also the ROD-$L^{2}$ method, for very high order polynomials, shows some parasitic modes that need higher CFL values to be stabilized.
@@ 7 Conclusions
Herein, we presented a thorough stability analysis for DG methods coupled with arbitrary high order embedded boundary treatments, in the context of hyperbolic equations. The numerical study was performed by visualizing the eigenspectrum of the discretized operators for a simplified system, and has provided critical insights into the stability properties of the considered geometrically unfitted approaches. To perform this study, it was necessary to write all embedded methods in matrix form, to allows for a straightforward integration into the global discretized operators. In this paper, we show for the first time that the ROD method can be recast as a polynomial correction that only depends on the basis function values on the real boundary and on the computational one, similarly to the one obtained for the SB method [ciallella2023shifted]. For all considered embedded treatments, it is interesting to notice the stability region is not symmetric when considering external ($d<0$) or internal ($d>0$) embedded boundaries. The findings of the analysis show that the ROD polynomial correction, when discarding the element crossed by the embedded boundary (i.e. for $d<0$ configurations), does not introduce any restriction on the CFL for polynomial degree $p\leq 3$ for the ROD-E version, and $p\leq 4$ for the ROD-$L^{2}$ version. This is an advantage with respect to the SB method, which shows a restriction on the CFL that becomes stricter as the polynomial degree increases. Instead, for $d>0$, all considered methods become unconditionally unstable beyond a critical distance $d>d_{max}(p)$, where $d_{max}$ decreases when increasing the polynomial degree. In the second part of the analysis, we study the impact of implicit time integration to understand whether it is possible to overcome the instabilities that arise with explicit methods. In all cases, we observe improvements in the stability properties of the embedded methods with implicit time integration. Moreover, for all configurations it is possible to obtain an unconditionally stable scheme. However for very high order methods, several configurations present lower bound conditions on the CFL: i.e. the CFL number should be high enough to stabilize the method. This is probably related to the fact that, even for implicit time integration, there are still some parasitic modes that make the method unstable and need larger CFL values to be stabilized. An extensive set of numerical experiments for the linear advection equation validate all results obtained by the stability analysis. The high order convergence slopes were recovered for all stable configurations $(d,\text{CFL})$, and instabilities were observed precisely where the analysis predicted them.
@@
Future work will extend this stability analysis to hyperbolic equations discretized by stabilized continuous Galerkin to investigate the impact of these embedded methods on a different finite element framework. Further implementations of the efficient ROD polynomial corrections in multiple dimensions will be pursued to improve computational performances on more realistic configurations. The possibility of perfoming a stability analysis based on the energy stability theory [carpenter1999stable, mattsson2003boundary] and on the G-K-S theory [gustafsson1972stability, strikwerda1980initial] will also be investigated.

### v2

@@ 6.2 ROD treatment and explicit time integration
The analysis for the ROD boundary treatment can be performed thanks to the reformulations carried out in section 5, which allows one to write easily ROD in matrix form.
@@
and the stability regions are given in figure 3.
The ROD-E method with polynomial degree $p=1,2,3$ shows stable areas for all values of $d\in[-1,0]$ under the standard CFL condition, an improvement over the SB method results in [6]. Instead, for $d\in[0,1]$, the method becomes unstable after a certain threshold $d>d_{max}$, which depends on the polynomial degree used to build the scheme, with the same $d_{max}$ observed for the SB method. Although the ROD-E method shows better stability properties for $p=1,2,3$ and distance $d\in[-1,0]$ than the SB method, for higher order polynomials $p=4,5,6$, also the ROD-E method becomes unconditionally unstable after a certain threshold. Contrary to the SB method, the ROD-E method never shows a constraint on the CFL number, it is either stable under a classical CFL condition or it becomes unconditionally unstable after a certain distance threshold.
For the $\mathbb{P}^{1}$ method, we can analytically compute the eigenvalues of the ROD-E boundary discretized operator, which read
@@
The scheme is stable in the semi-discrete sense when $d<2/3\approx 0.66$, which is the distance threshold after which the scheme becomes unconditionally unstable.
@@ 6.3 ROD treatment and implicit time integration
If we consider the implicit Euler time integration of the discontinuous Galerkin method with the ROD (either ROD-E or ROD-$L^{2}$) polynomial correction, the global system reads
@@ 7 Numerical experiments
For the numerical experiments, we consider equation (4) with a stationary manufactured solution
$u_{ex}(x)=0.1\sin(\pi x).$
$u_{ex}(x)$ is taken as initial condition for the numerical simulation. This makes the solution independent from the accuracy of time integration. The source term can be directly computed as
$\partial_{t}u_{ex}(x)+\partial_{x}u_{ex}(x)=s(x),\quad\text{resulting in}\quad s(x)=0.1\pi\cos(\pi x).$
@@
The one-dimensional domain is taken $\Omega=[0,2]$ with a Dirichlet boundary condition imposed at $x_{L}$. Since we are dealing with linear advection equation with positive speed, nothing is imposed at $x_{R}$. A set of uniform meshes with $N_{e}=20,40,80,160$ elements is considered for $\mathbb{P}^{p}$ with $p=1,2,3$. While another set of uniform meshes with $N_{e}=5,10,20,40$ elements is considered for $\mathbb{P}^{p}$ with $p=4,5,6$.
@@
The ROD (both ROD-E and ROD-$L^{2}$) polynomial corrections are imposed considering different distance values $d$ of the boundary from the left interface of the first cell. We refer to the previous analysis of the SB method in [6] to highlight the differences in stability properties between these two families of methods. When $d=0$, the boundary coincides with the left interface of the first cell, and therefore the correction is not applied. This is the case of standard fitted boundary conditions. Since all studied methods exhibit more instability when $d\in[0,\Delta x]$, the distances considered in the numerical experiments will be taken as $d\in[-\Delta x,0]$, meaning with the real boundary point located outside the first mesh element. For simplicity, in convergence tables, the distance is reported in non-dimensional form, i.e. $d/\Delta x$. Stability results obtained through the analysis are perfectly corroborated by the numerical results for all configurations of $d$, and for both explicit and implicit time integration.
@@
In table 1, we present the numerical results obtained with the ROD-E correction and explicit time integration. Again, we consider the greater distance that makes the scheme stable under the associated maximum CFL value. As also mentioned above, for $d<0$, it is possible to reach the maximum distance of d=1d=-1 for polynomials $p=1,2,3$ with no condition on the CFL. For $p=4,5,6$, the scheme is unconditionally unstable for d=1d=-1, and and the maximum stable distance is much lower: d=0.10d=-0.10 for $\mathbb{P}^{4}$, d=0.04d=-0.04 for $\mathbb{P}^{5}$, and d=0.015d=-0.015 for $\mathbb{P}^{6}$. Contrary to the SB method [6], there are no cases where it is possible to stabilize the ROD-E method by lowering the CFL. ROD-E is either stable under a classical CFL condition or unconditionally unstable; this is what happens for higher order polynomials after a certain distance threshold.
@@
In table 2, we present the numerical results obtained with the ROD-$L^{2}$ correction and explicit time integration. As before, we consider the greater distance that makes the scheme stable under the associated maximum CFL value. As also mentioned above, for $d<0$, it is possible reach the maximum distance of d=1d=-1 for polynomials $p=1,2,3,4$ with no condition on the CFL, meaning the $\mathbb{P}^{4}$ method is stable, unlike the ROD-E case. For $p=5,6$, the scheme is unconditionally unstable for d=1d=-1, and the maximum distance value for such polynomials is much lower: d=0.25d=-0.25 for $\mathbb{P}^{5}$, and d=0.05d=-0.05 for $\mathbb{P}^{6}$. Also in this case, there are no cases where it is possible to stabilize the ROD-$L^{2}$ method by lowering the CFL. ROD-$L^{2}$ is either stable under a classical CFL condition or unconditionally unstable. Similarly to what was observed for ROD-E.
@@
In table 3, we present the numerical results obtained with the ROD-E correction and implicit time integration. For implicit time integration, it is always possible to obtain a stable scheme for d=1d=-1 but for some cases the considered method has a lower bound condition on the CFL. For the ROD-E method with $p=1,2,3$ there are no conditions on the CFL, and it is therefore possible to use any CFL value and the scheme remains stable. This was expected since these schemes were also stable under the standard CFL condition with explicit time integration. In this case, it is interesting to see if coupling higher order methods with implicit time integration can result in stable schemes. However, for $p=4,5,6$ a lower bound on the CFL is necessary to ensure stability: for $p=4$ it is necessary to have $\text{CFL}\geq 3$, for $p=5$ we need $\text{CFL}\geq 6$, while for $p=6$ we have $\text{CFL}\geq 9$.
@@
In table 4, we present the numerical results obtained with the ROD-$L^{2}$ correction and implicit time integration. Also in this case, it is always possible to obtain a stable scheme for d=1d=-1. However, in some cases, the method has a lower bound condition on the CFL. For the ROD-$L^{2}$ method with $p=1,2,3,4$ there are no restruction on the CFL, and the scheme will stay stable. This was expected since these schemes were also stable under standard CFL condition with explicit time integration. Once again, we study if coupling higher order methods with implicit time integration can improve the stability for $p=5,6$. We show that for $p=5$ it is necessary to have $\text{CFL}\geq 0.7$, while for $p=6$ we need $\text{CFL}\geq 2$. Although less severe than for ROD-E, the ROD-$L^{2}$ method also exhibits parasitic modes for very high order polynomials that require higher CFL values for stabilization.
@@ 8 Conclusions
Herein, we presented a thorough stability analysis for DG methods coupled with arbitrary high order ROD embedded treatment, in the context of hyperbolic equations. The numerical study was performed by visualizing the eigenspectrum of the discretized operators for a simplified system, and has provided critical insights into the stability properties of the considered geometrically unfitted approaches. To perform this study, it was necessary to write all embedded methods in matrix form, to allows for a straightforward integration into the global discretized operators. In this paper, we show for the first time that the ROD method can be recast as a polynomial correction that only depends on the basis function values on the real boundary and on the computational one, similarly to the one obtained for the SB method in [8]. For all ROD versions, it is interesting to notice the stability region is not symmetric when considering external ($d<0$) or internal ($d>0$) embedded boundaries. The findings of the analysis show that the ROD polynomial correction, when discarding the element crossed by the embedded boundary (i.e. for $d<0$ configurations), does not introduce any restriction on the CFL for polynomial degree $p\leq 3$ for the ROD-E version, and $p\leq 4$ for the ROD-$L^{2}$ version. This is an advantage with respect to the SB method, which shows a restriction on the CFL that becomes stricter as the polynomial degree increases [6]. Instead, for $d>0$, all considered methods become unconditionally unstable beyond a critical distance $d>d_{max}(p)$, where $d_{max}$ decreases when increasing the polynomial degree. In the second part of the analysis, we study the impact of implicit time integration to understand whether it is possible to overcome the instabilities that arise with explicit methods. In all cases, we observe improvements in the stability properties of the ROD methods coupled with implicit time integration, and for all configurations it is possible to obtain an unconditionally stable scheme. However for very high order methods, several configurations present lower bound conditions on the CFL: i.e. the CFL number should be high enough to stabilize the method. This is probably related to the fact that, even for implicit time integration, there are still some parasitic modes that make the method unstable and need larger CFL values to be stabilized. An extensive set of numerical experiments for the linear advection equation validate all results obtained by the stability analysis. The high order convergence slopes were recovered for all stable configurations $(d,\text{CFL})$, and instabilities were observed precisely where the analysis predicted them.
@@
Future work will extend this stability analysis to hyperbolic equations discretized by stabilized continuous Galerkin to investigate the impact of these embedded methods on a different finite element framework. Further implementations of the efficient ROD polynomial corrections in multiple dimensions will be pursued to improve computational performances on more realistic configurations. The possibility of perfoming a stability analysis based on the energy stability theory [5, 21] and on the G-K-S theory [19, 30] will also be investigated.